A 13 ft board is cut into three pieces , two are equal to each other the third is 9inches shorter than the other two. what is the lengths of all boards?
Answer by macston(5194)

x=boards of equal length; x-9in=shorter board
x+x+x-9in=13ft convert ft to in
3x-9in=156in Add 9in to each side.
3x=165in
x=55in ANSWER 1: The boards of equal length are 55 inches.
x-9in=55in-9in=46in ANSWER 2: The short board is 46 inches.
CHECK:
55in+55in+46in=156in
156in=156in
13ft=13ft